OpenStudy (anonymous):

What is 12% written as a fraction? @shamim @Luigi0210 @jinxhead20 @adrynicoleb @elementwielder @dumbcow

OpenStudy (anonymous):

\[\frac{ 12 }{ 100 }\]

OpenStudy (anonymous):

that is the original fraction,..

OpenStudy (anonymous):

now simplify it to it lowest fraction

OpenStudy (anonymous):

*the

OpenStudy (anonymous):

\[\frac{ 6 }{ 50 }\] = ????

OpenStudy (anonymous):

3/25?

OpenStudy (anonymous):

I think so...YEP

OpenStudy (anonymous):

How do you make those fraction thingies with your keyboard???

OpenStudy (anonymous):

Just select the "equation" button and edit your equation.

OpenStudy (anonymous):

\[\frac{ 3 }{ 25 }\]

OpenStudy (anonymous):

yep lol